A firm produces 200 units hourly with $150 in labor costs an hour, $80 in materials costs an hour, and a fixed cost that works out to $20 per hour.

First, what are this firm's hourly variable and total costs?

a. VC=$1.15, TC=$1.25

b. VC=$150, TC=$170

c. VC=$170, TC=$250

d. VC=$230, TC=$250
Answer #1

Answer

Option d

d. VC=$230, TC=$250

Total cost =fixed cost +variable cost

variable cost =labor cost +material cost

=150+80

=$230

TC=20+230=$250